OpenStudy (anonymous):

A rectangle has a perimeter of 48 feet and a length of 14 feet. Which equation can you solve to find the width? 2n + 28 = 48 48 – n = 28 14 + n = 48 2n + 14 = 48

OpenStudy (anonymous):

im looking for short cut answers

OpenStudy (priyar):

what is the perimeter of the rectangle?

OpenStudy (anonymous):

it doesnt say

OpenStudy (priyar):

no..i am asking in general how do u find the perimeter of any rectangle?

OpenStudy (anonymous):

It does say what the perimeter is

OpenStudy (anonymous):

the perimeter says 48

OpenStudy (anonymous):

And how do you find the perimiter

OpenStudy (priyar):

in general, perimeter of a rectangle=2(l+b) l-length b- breadth

OpenStudy (priyar):

here width is "n".. so we have : perimeter=2(l+n)

OpenStudy (anonymous):

So in a more simple way, length + width. Width = n in this question. That's all the info you need to know for the answer

OpenStudy (priyar):

so 48=2l+2n.. now plug in the given length

OpenStudy (priyar):

@Redicle i think u mean 2(length + width)..!

OpenStudy (anonymous):

Oops, yeah

OpenStudy (priyar):

@kcunnin what do u get?

OpenStudy (kayders1997):

Okay so you have l is 14 plug that into your equation the equation is?

OpenStudy (megwatskies):

You would use 2n + 28 = 48 to find the width